# GRU: Gated Recurrent Unit with Spatial Attention for Spatio-Temporal Disaggregation

### v3

@@ Abstract.
Open data is frequently released spatially aggregated, usually to comply with privacy policies. But coarse, heterogeneous aggregations complicate learning and integration for downstream AI/ML systems. In this work, we consider models to disaggregate spatio-temporal data from a low-resolution, irregular partition (e.g., census tract) to a high-resolution, irregular partition (e.g., city block). We propose a model, Gated Recurrent Unit with Spatial Attention (GRU), where spatial attention layers are integrated into the original Gated Recurrent Unit (GRU) model. The spatial attention layers capture spatial interactions among regions, while the gated recurrent module captures the temporal dependencies. Additionally, we utilize containment relationships between different geographic levels (e.g., when a given city block is wholly contained in a given census tract) to constrain the spatial attention layers. For situations where limited historical training data is available, we study transfer learning scenarios and show that a model pre-trained on one city variable can be fine-tuned for another city variable using only a few hundred samples. Evaluating these techniques on two mobility datasets, we find that GRU provides a significant improvement over other neural models as well as typical heuristic methods, allowing us to synthesize realistic point data over small regions useful for training downstream models.
@@ 1. Introduction
High-quality, longitudinal, and freely available urban data, coupled with advances in machine learning, improve our understanding and management of urban environments (Han and Howe, 2023). Over the last two decades, cities have increasingly released datasets publicly on the web, proactively, in response to transparency regulation. For example, in the US, all 50 states and the District of Columbia have passed some version of the federal Freedom of Information (FOI) Act. While this first wave of open data was driven by FOI laws and made national government data available primarily to journalists, lawyers, and activists, a second wave of open data, enabled by the advent of open source and web 2.0 technologies, was characterized by an attempt to make data āopen by defaultā to civic technologists, government agencies, and corporationsĀ (Verhulst etĀ al., 2020).
@@
Open data affords study of urban dynamicsĀ (Vogel etĀ al., 2011; Yu etĀ al., 2017b), but, with few exceptionsĀ (Zhang etĀ al., 2020; Singh and Chen, 2023), models assume access to individual-level data, while privacy concerns motivate the release of aggregated dataĀ (Green etĀ al., 2017) such that aggregated data tends to be the norm in some fieldsĀ (Burrell etĀ al., 2004; Lozano etĀ al., 2009; Law etĀ al., 2018). Although aggregation can protect individual privacyĀ (Ullah etĀ al., 2021; Erkin etĀ al., 2013; He etĀ al., 2007) (without strong guaranteesĀ (Jia etĀ al., 2022; Brauneck etĀ al., 2023)), it complicates integration of multiple datasets (e.g., aligning tract-level to block-level data) and destroys information useful for training. For example, demographers have studied disaggregation techniques to uncover local hotspots erased by tract-level aggregation(Monteiro etĀ al., 2018). The New York City (NYC) taxi open datasets serve as another notable example, widely used in spatio-temporal traffic forecasting literature. Much of the work has relied on NYC taxi data collected before June 2016, when individual-level records were accessible (Mo and Gong, 2022; Zhang etĀ al., 2023; Jin etĀ al., 2022; Han and Howe, 2023). Subsequently, taxi records have been released at aggregated taxi zone levels, forcing applications to rely on out-of-date pre-2016 data.
@@
Motivated by these issues, we study the spatio-temporal disaggregation problem: to break down the aggregated urban data from a coarse, irregular spatial partitioning to a finer, irregular partitioning. The disaggregation problem has been studied in different domains. Spatial disaggregation methods in the demography literature tend to rely on heuristics: areal weighting, which assumes all variables are constant per unit areaĀ (Comber and Zeng, 2019), and dasymetric mappingĀ (Comber and Zeng, 2019), which assumes sufficient auxiliary data is available to predict the target variable using linear methods. Disaggregation has also been considered in the machine learning literature, outside of the spatiotemporal context, where the goal is to learn individual-level models given group-level aggregated dataĀ (Zhang etĀ al., 2020; Monteiro etĀ al., 2019a). However, these approaches explicitly seek a model where the target variable at the individual level is determined only by its parent groupās features as opposed to learning from global patterns across groups. In a spatiotemporal context, any region may influence any other: traffic on the bridge can increase wait times in midtown, for instance.
@@
In this paper, we evaluate neural architectures for the disaggregation problem. We proposed an overarching model, Gated Recurrent Unit with Spatial Attention (GRU), where spatial attention layers are incorporated in the Gated Recurrent Unit (GRU) model. The spatial attention layers capture the convoluted spatial interactions among regions, while the gated recurrent module captures the temporal dependencies. Specifically, the spatial attention layers calculate attention scores between the coarse and fine geographic regions. To enhance learning efficiency and effectiveness, we use (but do not necessarily require, containment: each fine level region is entirely contained in a coarse level region. Such phenomenon is practically seen in political boundaries. Exploiting this property, we create binary containment maps to be used in the spatial attention layers, promoting the generation of coherent and consistent values across varying geographic resolutions. We conduct three experiments targeting at different disaggregation levels. Our results show that GRU significantly enhances disaggregation performance and exhibits faster convergence compared to other neural models.
@@
In addition to our primary disaggregation findings, we conduct two experiments to enhance the practicality and the utility of the disaggregation task: 1) Transfer Learning. We relax the assumption that temporal historical data is always available and consider pre-training on variable A then fine-tuning on variable B. Our results demonstrate the remarkable effectiveness of transfer learning, yielding competitive outcomes with only a minimal number of data points. 2) Individual Points Synthesis. We first quantitatively show that our disaggregated values have a similar distribution to the ground truth values over a set of small block regions, via mutual information and KullbackāLeibler (KL) divergence statistics. Then we qualitatively demonstrate that a simple sampling scheme can be used to generate a realistic individual-level of points. The synthesized individual data can be further processed and used in other downstream tasks. Overall, we make the following contributions:
ā¢
We propose an original model, GRU, designed to disaggregate spatio-temporal data. The spatial attention layers model the spatial interactions while the gated recurrent module captures the temporal dependencies. By employing containment maps, our model is guided to produce coherent and consistent values across different geographic resolutions. On average, our model deliver an 2.2% improvement over the best baseline neural models (GRUs) and an 19.3% improvement over the best classical methods (historical ratio models). Additionally, GRU converge faster than other neural models.
ā¢
We evaluate the proposed model on two real mobility datasets addressing the scenario of the same city, but different variables. Our experiments demonstrate that the proposed model can enhance performance over the best baseline neural models (GRUs) on both datasets (2.3% on taxi data and 2.1% on bikeshare data).
ā¢
@@
We consider a transfer learning setting, under the assumption that sufficient training data is not always available to train a good neural model for every city variable. We show that in the same city, parameters learned from one variable can be used to make predictions on another variable, through fast fine-tuning with limited resources (12 hours v.s. 12 hours of pre-training), with competitive or sometimes even better performances.
ā¢
@@
After disaggregating to small regions, we synthesize individual-level data via sampling. We show that synthesized points have similar distributions to the true values, evidenced by high mutual information (0.736) and low KL-divergence (0.058) statistics.
@@ 2. Related Work
Spatial Disaggregation
@@
Spatial disaggregation is frequently used in demography to infer individual records from spatially aggregated regions such as census dataĀ (Qiu etĀ al., 2019; Wardropa etĀ al., 2018), in applications including disease mapping (Arambepola etĀ al., 2020) and vaccination coverage (Utazi etĀ al., 2018). A popular heuristic is areal weighting (AW), which assumes the value per unit area is constantĀ (Goodchild etĀ al., 1993). This method is common in practice due to its simplicity, general reliability, and ability to computed from a single variable. Pycnophylactic interpolation (PI), first introduced by Tobler in 1979 (Tobler, 1979), works similarly to areal weighting, but with a refinement to avoid discontinuities at the boundaries. PI distributes the values from source regions to target regions in an iterative manner, considering a weighted average of its nearest neighbors to avoid discontinuities while preserving the total count. Dasymetric mapping/interpolation/weighting (DM) approach is a family of methods that use the relationships between multiple variables to improve the estimated distribution. Unlike areal weighting, dasymetric mapping uses auxiliary information to generate a weighted distribution of source values onto the target regions. Recent methods incorporate machine learning approaches to model the relationships between auxiliary data and source values and determine the weights for disaggregation (Qiu etĀ al., 2019; Monteiro etĀ al., 2018, 2019b). Sources of auxiliary information include satellite images (Stevens etĀ al., 2015), 3D building information (Qiu etĀ al., 2019), mobile phone usage data (Deville etĀ al., 2014), terrain elevation and human settlement (Monteiro etĀ al., 2019b), and other variables more broadly available at high spatial and temporal resolution. Despite the popularity of dasymetric mapping methods in spatial disaggregation, there are several potential limitations of them (Comber and Zeng, 2019). They are wholly dependent on auxiliary variables, which are not always available (e.g., remote sensing images or land type data in population disaggregation tasks). Additionally, the performance of the model depends on the correlation between auxiliary data and source values, which may not be sufficiently high due to lack of relevant datasets or noisy measurements.
Learning From Aggregated Observations
@@
In addition to the spatial disaggregation methods, it is also possible to learn individual level probabilistic models from aggregated data, without first decomposing aggregated data into individual instances. The problem is generally referred to as learning from aggregated observations, where supervision labels are given to sets of instances instead of individual instances, while the goal is still to make inferences on unseen individuals (Zhang etĀ al., 2020; Singh and Chen, 2023; Law etĀ al., 2018). Zhang et al. (Zhang etĀ al., 2020) present a general probabilistic framework capable of accommodating a variety of aggregate observations for various tasks, such as classification and regression. They show that simple maximum likelihood solutions can be applied to various differentiable models such as deep neural networks and gradient boosting machines. Ma et al. (Ma etĀ al., 2020) studied the problem of learning nonlinear dynamics from aggregated data, where individual trajectory data is not available. They proposed a partial differential equation to describe the density evolution of aggregated data, which is subsequently combined with a Wasserstein generative adversarial network (WGAN) in the training process. Wei et al. (Wei etĀ al., 2023) studies a more specific problem ā classification from aggregate observations (CFAO), where the supervision is provided to groups of instances, instead of individual instances. They presented a novel universal method of CFAO, which generates unbiased estimators of the classification risk for arbitrary losses.
Spatio-Temporal Traffic Prediction
@@
Spatiotemporal disaggregation can be cast as a prediction problem, for which neural methods have received significant attention. Traffic forecasting is a spatiotemporal prediction problem that is a building block for a number of mobility applications(Jin etĀ al., 2023), such as trip planning (e.g., travel time estimation (Li etĀ al., 2018), route representation(Liu etĀ al., 2023b)) and resource optimization (e.g., dynamic fleet management (Zheng etĀ al., 2022)). Earlier workĀ (Li etĀ al., 2017; Li and Zhu, 2021; Yu etĀ al., 2017a) used graph convolutional network (GCNs) to capture spatial interactions and sequential models such as Long-Short Term Memory (LSTMs) or Gated Recurrent Unit (GRUs), to model temporal dependencies. Later work (Zhu etĀ al., 2021; Zheng etĀ al., 2020; Lan etĀ al., 2022) utilized spatial and/or temporal attention mechanisms to capture more complicated spatial and temporal dynamics and correlations. Recent work applies spatio-temporal transformers(Yan etĀ al., 2021; Liu etĀ al., 2023a; Fang etĀ al., 2022). Spatio-temporal transformers have also been applied in other areas, such as visual tracking (Yan etĀ al., 2021), trajectory prediction(Chen etĀ al., 2021), and human action recognition(Ahn etĀ al., 2023).
@@ 3. Datasets
In our research, we use two mobility datasets ā NYC taxi data and NYC bikeshare data. Each dataset contains individual-level records.
ā¢
@@
ā¢
NYC Bikeshare Data: NYC bikeshare data is collected from NYC DOT 333https://citibikenyc.com/system-datafrom 01/01/2021 to 06/30/2021. Similar to the taxi data, the raw data are presented in tabular format. Each data point represents one bike trip, including the longitude and latitude of the location where the bike was unlocked.
These two datasets represent two examples of mobility behavior in the same built environment, affording comparisons under similar conditions. Moreover, these datasets afford transfer learning studies (e.g., training on taxi data and evaluating on bikeshare data) to address the situation where available data may be limited.
We have restricted our experiments to the Manhattan region. We use four distinct pre-defined geographic levels offering varying levels of spatial resolution: Public Use Microdata Areas (PUMA), Neighborhood Tabulation Areas (NTA), Census Tracts (TRACT), and Census Blocks (BLOCK) (see Figure 1). We obtain the 2010 geographic boundaries for these levels from the NYC open data portal. Figure 1 shows the resolutions of four geographic levels.

### v4

# SARN: Structurally-Aware Recurrent Network for Spatio-Temporal Disaggregation

@@ Abstract.
Open data is frequently released spatially aggregated, usually to comply with privacy policies. But coarse, heterogeneous aggregations complicate learning and integration for downstream AI/ML systems. In this work, we consider models to disaggregate spatio-temporal data from a low-resolution, irregular partition (e.g., census tract) to a high-resolution, irregular partition (e.g., city block). We propose an overarching model named the Structurally-Aware Recurrent Network (SARN), which integrates structurally-aware spatial attention (SASA) layers into the Gated Recurrent Unit (GRU) model. The spatial attention layers capture spatial interactions among regions, while the gated recurrent module captures the temporal dependencies. Each SASA layer calculates both global and structural attention — global attention facilitates comprehensive interactions between different geographic levels, while structural attention leverages the containment relationship between different geographic levels (e.g., a city block being wholly contained within a census tract) to ensure coherent and consistent results. For scenarios with limited historical training data, we explore transfer learning and show that a model pre-trained on one city variable can be fine-tuned for another city variable using only a few hundred samples. Evaluating these techniques on two mobility datasets, we find that on both datasets, SARN significantly outperforms other neural models (5% and 1%) and typical heuristic methods (40% and 14%), enabling us to generate realistic, high-quality fine-grained data for downstream applications.
@@ 1. Introduction
High-quality, longitudinal, and freely available urban data, coupled with advances in machine learning, improve our understanding and management of urban environments (Han and Howe, 2023). Over the last two decades, cities have increasingly released datasets publicly on the web, proactively, in response to transparency regulation. For example, in the US, all 50 states and the District of Columbia have passed some version of the federal Freedom of Information (FOI) Act. While this first wave of open data was driven by FOI laws and made national government data available primarily to journalists, lawyers, and activists, a second wave of open data, enabled by the advent of open source and web 2.0 technologies, was characterized by an attempt to make data “open by default” to civic technologists, government agencies, and corporations (Verhulst et al., 2020).
@@
Open data affords study of urban dynamics (Vogel et al., 2011; Yu et al., 2017b), but, with few exceptions (Zhang et al., 2020; Singh and Chen, 2023), models assume access to individual-level data, while privacy concerns motivate the release of aggregated data (Green et al., 2017) such that aggregated data tends to be the norm in some fields (Burrell et al., 2004; Lozano et al., 2009; Law et al., 2018). Although aggregation can protect individual privacy (Ullah et al., 2021; Erkin et al., 2013; He et al., 2007) (without strong guarantees (Jia et al., 2022; Brauneck et al., 2023)), it complicates integration of multiple datasets (e.g., aligning tract-level to block-level data) and destroys information useful for training. For example, demographers have studied disaggregation techniques to uncover local hotspots erased by tract-level aggregation(Monteiro et al., 2018). The New York City (NYC) taxi open datasets serve as another notable example, widely used in spatio-temporal traffic forecasting literature. Much of the work has relied on NYC taxi data collected before June 2016, when individual-level records were accessible (Mo and Gong, 2022; Zhang et al., 2023; Jin et al., 2022; Han and Howe, 2023). Subsequently, taxi records have been released at aggregated taxi zone levels, forcing applications to rely on out-of-date pre-2016 data.
@@
Motivated by these issues, we study the spatio-temporal disaggregation problem — to break down the aggregated urban data from a coarse, irregular spatial partitioning to a finer, irregular partitioning. The disaggregation problem has been studied in different domains. Spatial disaggregation methods in the demography literature tend to rely on heuristics: areal weighting, which assumes all variables are constant per unit area (Comber and Zeng, 2019), and dasymetric mapping (Comber and Zeng, 2019), which assumes sufficient auxiliary data is available to predict the target variable using linear methods. Disaggregation has also been considered in the machine learning literature, outside of the spatiotemporal context, where the goal is to learn individual-level models given group-level aggregated data (Zhang et al., 2020; Monteiro et al., 2019a). However, these approaches explicitly seek a model where the target variable at the individual level is determined only by its parent group’s features as opposed to learning from global patterns across groups. In a spatiotemporal context, any region may influence any other: traffic on the bridge can increase wait times in midtown, for instance.
@@
In this paper, we evaluate neural architectures for the disaggregation problem, and demonstrate the utility of the disaggregated data (See Figure 1). We proposed an overarching model, Structurally-Aware Recurrent Network (SARN), which integrates structurally-aware spatial attention layers (SASA) into the Gated Recurrent Unit (GRU) model. The SASA layers capture convoluted spatial interactions among regions, while the gated recurrent module captures the temporal dependencies. Specifically, each SASA layer calculates two types of attention — global and structural. Global attention attends each coarse region to all fine regions, facilitating a comprehensive, macro-level geographic interactions. To enhance learning efficiency and effectiveness, we leverage the structural property exhibiting between geographic levels, containment — each fine level region is entirely contained in a coarse level region. Such phenomenon is practically seen in political boundaries. Exploiting this structural property, we create binary containment maps to be used in structural attention calculation, promoting the generation of coherent and consistent values across varying geographic levels. We conduct three experiments targeting various disaggregation levels with two mobility datasets from NYC. Our results demonstrate that SARN significantly enhances disaggregation performance and converges faster than other neural models.
@@
In addition to our primary disaggregation findings, we conduct three supplementary experiments to enhance the practicality of the disaggregation task and demonstrate the utility of our model: 1) Transfer Learning. We relax the assumption that temporal historical data is always available and explore pre-training on variable , then fine-tuning on variable . Our results demonstrate the remarkable effectiveness of transfer learning, yielding competitive outcomes compared with specialized models, with only a minimal number of samples. 2) Downstream Application. The disaggregated data can be valuable for downstream applications. We use traffic demand prediction as an example, as accurate traffic demand prediction is essential for cities to allocate transportation resources effectively, especially in fine regions (e.g., city blocks), potentially reducing traffic congestion. We train several models with the disaggregated taxi data and compare their performance to those trained with ground truth data. Our results show remarkable authenticity of the disaggregated data compared to the real data. 3) Individual Points Synthesis. We quantitatively show that our disaggregated values have a similar distribution to the ground truth values over a set of regions, via mutual information and Kullback–Leibler divergence statistics. Then we qualitatively demonstrate that a simple sampling scheme can be used to generate a realistic individual-level of points.
Overall, we make the following contributions:
We propose the Structurally-Aware Recurrent Network (SARN), designed to disaggregate spatio-temporal data. The structurally-aware spatial attention (SASA) layers model spatial interactions, while the gated recurrent module captures temporal dependencies. Within each SASA layer, global attention is calculated to enable macro-level interactions between levels. By employing containment maps, structural attention is calculated to produce coherent and consistent values between different levels.
We test three disaggregation resolutions and two mobility datasets. On NYC taxi data, SARN achieves a 5% improvement over the best neural models and a 40.1% improvement over the best classical methods on average. For NYC bikeshare data, the average improvements are 1.2% and 13.6%, respectively. Furthermore, SARN converges faster than other neural models.
@@
We consider a transfer learning scenario, acknowledging that sufficient training data may not always be available to train an effective neural model for every city variable. Our results demonstrate that, within the same city, parameters learned from one variable can be used to make predictions on another variable through rapid fine-tuning (1-2 hours versus 12 hours of pre-training), achieving competitive performance.
@@
We demonstrate the utility of disaggregated data in downstream traffic demand prediction. The prediction difference between models trained with disaggregated values and ground truth values is trivial (0.78%), indicating that the disaggregated data is of high-quality and closely matches the ground truth values. Additionally, we synthesize individual-level data via sampling and show that synthesized points have similar distributions to the true values, evidenced by high mutual information (0.736) and low KL-divergence (0.058) statistics.
@@ 2. Related Work
2.1. Spatial Disaggregation
Spatial disaggregation is frequently used in demography to infer individual records from spatially aggregated regions such as census data (Qiu et al., 2019; Wardropa et al., 2018), frequently applied in population analysis (e.g., dynamic population mapping). Dynamic data on human population distribution at high resolution is important for a wide spectrum of activities and real-life applications, such as urban applications of business locations, transportation planning, city service management etc. (Zong et al., 2019) Other applications include disease mapping (Arambepola et al., 2020) and vaccination coverage (Utazi et al., 2018).
A popular heuristic is areal weighting (AW), which assumes the value per unit area is constant (Goodchild et al., 1993). This method is common in practice due to its simplicity, general reliability, and ability to computed from a single variable. Pycnophylactic interpolation (PI), first introduced by Tobler in 1979 (Tobler, 1979), works similarly to areal weighting, but with a refinement to avoid discontinuities at the boundaries. PI distributes the values from source regions to target regions iteratively , considering a weighted average of its nearest neighbors to avoid discontinuities while preserving the total count.
@@ 2.1. Spatial Disaggregation
Dasymetric mapping/interpolation/weighting (DM) approach is a family of methods that use the relationships between multiple variables to improve the estimated distribution. Unlike areal weighting, dasymetric mapping uses auxiliary information to generate a weighted distribution of source values onto the target regions. Recent methods incorporate machine learning approaches to model the relationships between auxiliary data and source values and determine the weights for disaggregation (Qiu et al., 2019; Monteiro et al., 2018, 2019b). Sources of auxiliary information include satellite images (Stevens et al., 2015), 3D building information (Qiu et al., 2019), mobile phone usage data (Deville et al., 2014), terrain elevation and human settlement (Monteiro et al., 2019b), and other variables more broadly available at high spatial and temporal resolution. Despite the popularity of dasymetric mapping methods in spatial disaggregation, there are several potential limitations of them (Comber and Zeng, 2019). They are wholly dependent on auxiliary variables, which are not always available (e.g., remote sensing images or land type data in population disaggregation tasks). Additionally, the performance of the model depends on the correlation between auxiliary data and source values, which may be low due to lack of relevant datasets or noisy measurements.
Deep learning methods have also been directly applied to the disaggregation problem in recent years. For example, Zong et al. (Zong et al., 2019) proposed the DeepDPM model, which uses a super-resolution CNN to analyze both spatial and temporal patterns using coarse data and point-of-interest information. In the urban computing domain, several recent studies have focused on generating fine-grained traffic flow from coarse-grained data using super-resolution techniques. Ouyang et al. (Ouyang et al., 2020) introduced the UrbanFM model, which employs a feature extraction module and a distributional upsampling module to infer fine-grained crowd flow from coarse observations. Liu et al. (Liu et al., 2023c) proposed the RATFM model, which leverages prior knowledge of road networks to learn the road-aware spatial distribution of fine-grained traffic flow. Both studies use super-resolution techniques, but their problem setting differs slightly from ours. Their coarse observations are at the pixel level, while our coarse values are at predefined geopolitical region levels (e.g., tracts, blocks). In an urban image, each region may cover numerous pixels that share the same value.
2.2. Learning From Aggregated Observations
@@
In addition to the spatial disaggregation methods, it is also possible to learn individual level probabilistic models from aggregated data, without first decomposing aggregated data into individual instances. The problem is generally referred to as learning from aggregated observations, where supervision labels are given to sets of instances instead of individual instances, while the goal is still to make inferences on unseen individuals (Zhang et al., 2020; Singh and Chen, 2023; Law et al., 2018). Zhang et al. (Zhang et al., 2020) present a general probabilistic framework capable of accommodating a variety of aggregate observations for various tasks, such as classification and regression. They show that simple maximum likelihood solutions can be applied to various differentiable models such as deep neural networks and gradient boosting machines. Ma et al. (Ma et al., 2020) studied the problem of learning nonlinear dynamics from aggregated data, where individual trajectory data is not available. They proposed a partial differential equation to describe the density evolution of aggregated data, which is subsequently combined with a Wasserstein generative adversarial network (WGAN) in the training process. Wei et al. (Wei et al., 2023) studies a more specific problem — classification from aggregate observations (CFAO), where the supervision is provided to groups of instances, instead of individual instances. They presented a novel universal method of CFAO, which generates unbiased estimators of the classification risk for arbitrary losses.
2.3. Spatio-Temporal Traffic Prediction
@@
Spatiotemporal disaggregation can be cast as a prediction problem, for which neural methods have received significant attention. Traffic forecasting is a spatiotemporal prediction problem that is a building block for a number of mobility applications(Jin et al., 2023), such as trip planning (e.g., travel time estimation (Li et al., 2018), route representation(Liu et al., 2023b)) and resource optimization (e.g., dynamic fleet management (Zheng et al., 2022)). Earlier work (Li et al., 2017; Li and Zhu, 2021; Yu et al., 2017a) used graph convolutional network (GCNs) to capture spatial interactions and sequential models such as Long-Short Term Memory (LSTMs) or Gated Recurrent Unit (GRUs), to model temporal dependencies. Later work (Zhu et al., 2021; Zheng et al., 2020; Lan et al., 2022) utilized spatial and/or temporal attention mechanisms to capture more complicated spatial and temporal dynamics and correlations. Recent work applies spatio-temporal transformers(Yan et al., 2021; Liu et al., 2023a; Fang et al., 2022). Spatio-temporal transformers have also been applied in other areas, such as visual tracking (Yan et al., 2021), trajectory prediction(Chen et al., 2021), and human action recognition(Ahn et al., 2023).
@@ 3. Datasets
In our research, we use two mobility datasets — NYC taxi data and NYC bikeshare data. Each dataset contains individual-level records.
@@
NYC Bikeshare Data: NYC bikeshare data is collected from NYC DOT 333https://citibikenyc.com/system-datafrom 01/01/2021 to 12/31/2021. Similar to the taxi data, the raw data are presented in tabular format. Each data point represents one bike trip, including the longitude and latitude of the location where the bike was unlocked. The dataset spans the entire year, allowing the inspection of seasonal variations.
These datasets represent two examples of mobility events in the same city (taxi and bikeshare in NYC). This collection enables study of variance of our method (same city, same application domain, but different variable). Moreover, NYC taxi and bikeshare datasets afford transfer learning studies (e.g., training on taxi data and evaluating on bikeshare data) to address the situation where available data may be limited. For NYC, we have restricted our experiments to the Manhattan region. We use four pre-defined geographic levels offering varying levels of spatial resolution: Public Use Microdata Areas (PUMA), Neighborhood Tabulation Areas (NTA), Census Tracts (TRACT), and Census Blocks (BLOCK) (see Figure 2). We obtain the 2010 boundaries for these levels from the NYC open data portal. Figure 2 shows the resolutions of four geographic levels.
Data | PUMA | NTA | TRACT | BLOCK |
NYC Taxi | 1443.01 | 450.94 | 51.17 | 3.86 |
(1776.03) | (689.16) | (73.32) | (8.03) | |
NYC Bikeshare | 189.24 | 59.14 | 6.71 | 0.50 |
(279.28) | (97.86) | (13.09) | (2.51) |